if a weak acid, HA, is 3% dissociated in a 0.25M solution, calculate the Ka and the pH of the solution.
1
Expert's answer
2012-02-17T08:34:14-0500
a=3% (0.03)
C=0.25M
Ka-? pH-?

HA <=>H+ +
A-

a=(K/C) ^ (1/2)
a^ 2= K/C
K=(a^ 2)/C
K= (0.03^ 2)
/0.25=0.0036

[H+]=(K*C) ^ (1/2)=0.03
pH=-log[H+]=1.52

K=0.0036
pH=1.52